OpenStudy (anonymous):

4^(2X+3)=1 Since the 4 and 1 don't have a common base I get a bit confused.

OpenStudy (anonymous):

Recall that \(4^0=1\). In fact, for \(a>0\), \(a^0=1\). So, you have \[4^{2x+3}=4^0\] which gives \[2x+3=0\]

OpenStudy (anonymous):

ahh thanks.